OpenStudy (anonymous):

A random sample of 1000 oranges showed that the mean amount of juice per orange was 7 fluid ounces, with a standard deviation of 1.6 fluid ounces. If the z-score for a particular orange was –1.5, how much juice was produced by this orange? Round approximate values to the nearest tenth of a fluid ounce.

OpenStudy (amistre64):

7 -1.5 = 5.5

OpenStudy (amistre64):

5.5\14 ?

OpenStudy (anonymous):

7-1.5*1.6=4.6

OpenStudy (anonymous):

well their is a 5.5 and 4.6

OpenStudy (amistre64):

5.5 is its position

OpenStudy (anonymous):

im looking for a calculator

OpenStudy (anonymous):

The z score is standarised and tells you how many standard deviations away from the mean the orange is.

OpenStudy (anonymous):

i really hate normal distribution

OpenStudy (anonymous):

its not so normal

OpenStudy (anonymous):

In a certain normal distribution of scores, the mean is 50 and the standard deviation is 4. Find the z-score corresponding to a score of 55.

OpenStudy (anonymous):

Yes, it is normal. This orange is 1.5 standard deviations below the mean and the mean is 7 so the orange produces 7-1.5*1.6=4.6 fluid onces.

OpenStudy (anonymous):

well im do this for the first time on the web... i should have done it in a class room with a real teacher

OpenStudy (amistre64):

i get the 5.5 .... but whats the logic behind *zscore?

OpenStudy (amistre64):

or is that 7-(1.5*1.6)

OpenStudy (anonymous):

\[Z=\frac{X-\mu}{\sigma}\] so it tells you how many standard deviations below the mean a certain value is.

OpenStudy (anonymous):

Yes, 7-(1.5*1.6)

OpenStudy (amistre64):

stats class is a little confusing to me becasue the teacher is just throwing formulas at us with no explanation ...

OpenStudy (anonymous):

50-(4*55)=

OpenStudy (anonymous):

In a certain normal distribution of scores, the mean is 50 and the standard deviation is 4. Find the z-score corresponding to a score of 55.

OpenStudy (anonymous):

did i do it right lol

OpenStudy (anonymous):

No, see the formula I posed earlier. It should be (55-50)/4

OpenStudy (anonymous):

what confuses me is the signs for the formula... but i got it

OpenStudy (anonymous):

this is my first time ever doing this plez excuse me

OpenStudy (anonymous):

Ah, ok, sorry. X is the number you're trying to find the Z value for, mu is the mean and sigma is the standard deviation.

OpenStudy (anonymous):

but i understand it better than what the book would explain... its like 10 pages explain that formula

OpenStudy (anonymous):

Remember, the z score just tells you how many standard deviations away from the mean a value is. So, say I had a mean of 50, and an SD of 2, if I was trying to find the z score of 56, it would be 3 because 56 is 3 standard deviations from the mean.